# Dyonic RN-like and Taub-NUT-like black holes in Einstein-bumblebee gravity

### v1

@@ III.1 Exact dyonic solution
where $d\Omega_{2,k}^{2}={du^{2}}/{(1-ku^{2})}+(1-ku^{2})d\varphi^{2}$ with k=1,0,1k=1,0,-1, corresponding to the metric of the unit 2-sphere, the 2-torus or the unit hyperbolic 2-space, respectively. The constant $p$ denotes the magnetic charge parameter. For the bumblebee field $B_{\mu}$, following Ref. [16], we consider a nonzero VEV $b_{\mu}$ oriented along a radial direction, taking the form
@@
$f=\frac{1}{1+b^{2}\gamma}\left[k-\frac{m}{r}+\frac{(1-2b^{2}(2\gamma_{1}+\gamma_{2}))q^{2}}{4r^{2}}+\frac{\left(1-4b^{2}\gamma_{1}\right)p^{2}}{4r^{2}}\right]\,$, (23)
@@
$h$ $=$ $k-\frac{m}{r}+\frac{q^{2}}{2\left(2+\ell\right)r^{2}}+\frac{\left(1+\ell\right)p^{2}}{2\left(2+3\ell\right)r^{2}}\,,\quad f=\frac{h}{1+\ell}\,,\quad\phi=\frac{q}{r}\,$, (26)
@@
Note that we have chosen the potential $\phi$ to vanish at infinity. In the limit $\ell\to 0$, the solution reduces to the standard dyonic RN black hole with a generic topological horizon. Alternatively, in the absence of both electric and magnetic charges with $k=1$, the solution reduces to a Schwarzschild-like black hole, which has been extensively discussed in Ref. [16]. The Kretschmann scalar, i.e., the square of the Riemann tensor, can be easily calculated as
@@
which exhibits the similar singularity structure as the dyonic RN case, but clearly deviates from the corresponding Kretschmann invariant when $\ell\neq 0$.
@@ III.2 The difference from [18]
$h$ $=$ $k-\frac{m}{r}+\frac{(1+\ell)q^{2}}{2\left(2+\ell\right)r^{2}}\,,\qquad f=\frac{h}{1+\ell}\,,\qquad\phi=\frac{\sqrt{\ell+1}q}{r}\,,\qquad b_{r}=\frac{b}{\sqrt{f}}\,.$ (30)
@@
$h$ $=$ $k-\frac{m}{r}+\frac{(1+\ell)q^{2}}{2\left(2+\ell\right)r^{2}}+\frac{(\ell+1)(3\ell+2)p^{2}}{2(\ell+2)^{2}r^{2}}\,,\qquad\gamma_{2}=-\frac{2\gamma(\ell+1)}{(\ell+2)^{2}}\frac{p^{2}}{q^{2}}\,.$ (32)
@@
$A_{{{\scriptscriptstyle(1)}}}$ $=$ $\phi(r)dt+pqud\varphi\,,\qquad h=k-\frac{m}{r}+\frac{(1+\ell)q^{2}}{2\left(2+\ell\right)r^{2}}+\frac{(\ell+1)(3\ell+2)p^{2}q^{2}}{2(\ell+2)^{2}r^{2}}\,$, (34)
@@ IV Thermodynamics
Next, in order to obtain the electric charge $Q_{e}$, it is convenient to introduce an antisymmetric tensor ${\cal F}^{\mu\nu}$
@@
The Wald entropy formula [94, 95] reads
@@ IV.1 Wald formalism
A direct consequence of the Wald formalism is the Wald entropy formula, which extends the Bekenstein-Hawking entropy and possesses remarkable universality. In higher-derivative gravity theories, black hole entropy is no longer proportional to the horizon area. The Wald entropy formula, however, correctly accounts for these corrections [106, 107, 108] and is thus widely used in many black hole–related effective theories [110, 104, 105, 109]. Thus far, the Wald formalism and the Wald entropy formula have been applied to a broad range of extended gravity theories, including tensor-scalar theories [111, 112, 102, 103], Maxwell theory and its extensions [113, 114, 115, 113, 106, 99, 100], as well as the neutral bumblebee theory [49, 19].
In fact, the Wald entropy formula is simply a corollary of the Wald formalism and can be obtained from it only under specific conditions. Indeed, previous studies of Horndeski theory and bumblebee theory have demonstrated that the Wald entropy formula fails to provide a black hole entropy consistent with the first law of thermodynamics [49, 19, 102, 103]. More precisely, the dependence of the Noether charge on the Killing vector involves only $\xi^{\mu}$ and $\nabla^{\mu}\xi^{\nu}$. On the black hole horizon, if all fields are smooth, one has $\xi\to 0$ and $\nabla^{\mu}\xi^{\nu}\to K\epsilon^{\mu\nu}$, from which the Wald entropy formula follows. However, the $B_{\mu}$ field (15, 16) diverges on the black hole horizon, rendering the Wald entropy formula inapplicable. Hence, the complete thermodynamics must be derived directly from the original Wald formalism.
@@
At horizon, we have
@@
Thus the first law of black hole thermodynamics reads
@@
It is evident that, although the Lorentz-violating parameter $\ell$ affects nearly all thermodynamic quantities of the dyonic RN-like black hole—except for the electric potential and magnetic charge—all forms of the first law of black hole thermodynamics, Eqs. (88)–(90), remain consistent with their GR counterparts.
@@ V.1 Exact solution
$ds^{2}$ $=$ $-h(r)(dt+2Nud\varphi)^{2}+\frac{dr^{2}}{f(r)}+(r^{2}+N^{2})\Big(\frac{du^{2}}{1-ku^{2}}+(1-ku^{2})d\varphi^{2}\Big)\,$, (91)
$A_{{{\scriptscriptstyle(1)}}}$ $=$ $\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Big[\phi(r)(dt+2Nud\varphi)+pud\varphi\Big]+c_{e}dt\,,\qquad B_{{{\scriptscriptstyle(1)}}}=b_{r}(r)dr\,$ (92)
where
$\phi(r)$ $=$ $\frac{1}{2N}\Bigg\{q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]+p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]-p\Bigg\}\,$, (93)
$b_{r}(r)$ $=$ $\frac{b}{\sqrt{f(r)}}\,,\qquad f(r)=\frac{h(r)}{1+\ell}\,,\qquad h(r)=\frac{2k(r^{2}-N^{2})-2mr+\frac{p^{2}+q^{2}}{\ell+2}}{2(r^{2}+N^{2})}\,,\qquad\ell=b^{2}\gamma\,.$ (94)
For $N=0$, the solution reduces to that presented in Section III. For convenience, we redefine the magnetic charge parameter $p$, which differs from the one in (14) by a constant factor $\sqrt{(3\ell+2)/((\ell+1)(\ell+2))}$. The parameter $c_{e}$ is a gauge constant, which does not affect the solution. In previous RN-like solution, it was set to zero.
@@ V.2 Normal thermodynamic quantities
$T=\frac{1}{4\pi\sqrt{\ell+1}}\frac{1}{r_{h}}\Big[1-\frac{p^{2}+q^{2}}{2(\ell+2)(r_{h}^{2}+N^{2})}\Big]\,.$ (95)
Using the expression for the Maxwell field $A_{{{\scriptscriptstyle(1)}}}$ and its EOM, one can determine the electric potential $\Phi_{e}$ and the electric charge $Q_{e}$. Likewise, the magnetic charge $Q_{m}$ (43) can be obtained via the Bianchi identity $dF_{{{\scriptscriptstyle(2)}}}=0$, and the magnetic potential $\Phi_{m}$ (45) can be calculated by introducing the dual 1-form field $\mathcal{A}_{{{\scriptscriptstyle(1)}}}$ (44)
@@
$\mathcal{A}_{t}$ $=$ $-\frac{\sqrt{\ell+1}}{2\kappa N(\ell+2)}\Bigg\{q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]$ (98)
$-p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]\Bigg\}\,$,
$\mathcal{A}_{\varphi}$ $=$ $-\frac{u\sqrt{\ell+1}}{\kappa(\ell+2)}\Bigg\{q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]$ (100)
$-p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]\Bigg\}\,.$
Here, $c_{m}$ is a gauge parameter, which, similarly to the RN-like case, we also set to zero. The electric and magnetic potentials (42,45) are given by
$\Phi_{e}$ $=$ $\frac{1}{2N}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Bigg\{-p+p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (102)
$+q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,$,
$\Phi_{m}$ $=$ $\frac{\sqrt{\ell+1}}{2\kappa N(\ell+2)}\Bigg\{q-q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (104)
$+p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,.$
Because of the Misner string, computing the electric charge requires careful treatment of the integration. Here,
$Q_{e}$ $=$ $\frac{1}{2\kappa}\int\mathcal{Q}_{tr}(r)dt\wedge dr+\mathcal{Q}_{r\varphi}(r,u)dr\wedge d\varphi+\mathcal{Q}_{u\varphi}(r)du\wedge d\varphi\,$, (105)
$\mathcal{Q}_{r\varphi}$ $=$ $-2Nu\mathcal{Q}_{tr}\,$, (106)
$\mathcal{Q}_{tr}$ $=$ $\frac{\ell+1}{\kappa\sqrt{(\ell+2)(3\ell+2)}(r^{2}+N^{2})}\Bigg\{p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (108)
$+q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,$,
$\mathcal{Q}_{u\varphi}$ $=$ $-\frac{\sqrt{\ell+1}}{\kappa(\ell+2)}\Bigg\{q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (110)
$-p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,.$
@@
$Q_{e}$ $=$ $\frac{1}{2\kappa}\int d\varphi\Big[\int_{-1}^{1}\mathcal{Q}_{u\varphi}(r\to\infty)du^{\prime}+\int_{r_{h}}^{r\to\infty}\mathcal{Q}_{r\varphi}(r^{\prime},u)dr^{\prime}\Big]=-\frac{w_{2}}{2}\mathcal{A}_{\varphi}|_{u=-1}^{u=1}$ (111)
$=$ $\frac{w_{2}\sqrt{\ell+1}}{\kappa(\ell+2)}\Bigg\{q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (113)
$-p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,$,
can be determined. The magnetic charge is calculated using the same integration procedure
$Q_{m}$ $=$ $\frac{1}{2\kappa}\int\mathcal{P}_{tr}(r)dt\wedge dr+\mathcal{P}_{r\varphi}(r,u)dr\wedge d\varphi+\mathcal{P}_{u\varphi}(r)du\wedge d\varphi\,$, (114)
$\mathcal{P}_{r\varphi}$ $=$ $-2Nu\mathcal{P}_{tr}\,$, (115)
$\mathcal{P}_{tr}$ $=$ $-\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\phi^{\prime}(r)\,,\qquad\mathcal{P}_{u\varphi}=\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\big(p+2N\phi(r)\big)\,$, (116)
resulting in $Q_{m}$ given by
$Q_{m}$ $=$ $\frac{1}{2\kappa}\int d\varphi\Big[\int_{-1}^{1}\mathcal{P}_{u\varphi}(r\to\infty)du^{\prime}+\int_{r_{h}}^{r\to\infty}\mathcal{P}_{r\varphi}(r^{\prime},u)dr^{\prime}\Big]=\frac{w_{2}}{2}A_{\varphi}|_{u=-1}^{u=1}$ (117)
$=$ $w_{2}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Bigg\{p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (119)
$+q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,.$

### v2

@@ III.1 Exact dyonic solution
where
$d\Omega_{2,k}^{2}=\frac{du^{2}}{1-ku^{2}}+(1-ku^{2})d\varphi^{2}\,$, (15)
with k=1,0,1k=1,0,-1, corresponding to the metric of the unit 2-sphere, the 2-torus or the unit hyperbolic 2-space, respectively. The constant $p$ denotes the magnetic charge parameter. For the bumblebee field $B_{\mu}$, following Ref. [16], we consider a nonzero VEV $b_{\mu}$ oriented along a radial direction, taking the form
@@
$f=\frac{1}{1+b^{2}\gamma}\Bigg[k-\frac{m}{r}+\frac{(1-2b^{2}(2\gamma_{1}+\gamma_{2}))q^{2}}{4r^{2}}+\frac{\left(1-4b^{2}\gamma_{1}\right)p^{2}}{4r^{2}}\Bigg]\,$, (24)
@@
$h$ $=$ $k-\frac{m}{r}+\frac{q^{2}}{2\left(2+\ell\right)r^{2}}+\frac{\left(1+\ell\right)p^{2}}{2\left(2+3\ell\right)r^{2}}\,$,
$f$ $=$ $\frac{h}{1+\ell}\,,\quad\phi=\frac{q}{r}\,$, (27)
@@
Note that we have chosen the potential $\phi$ to vanish at infinity. In the expression for the metric function $h$, the Lorentz-violating parameter $\ell$ does not appear explicitly in the mass term, since it can be absorbed into the mass integration constant $m$. However, once electromagnetic fields are coupled, this parameter can no longer be eliminated simultaneously from the metric function $h$ and $A_{\mu}$ through a redefinition of the electric and magnetic integration constants $q$ and $p$. More importantly, as in the neutral black-hole case, the Lorentz-violating parameter $\ell$ spoils the simple relation $f=h$. This reflects a fundamental feature of Einstein-bumblebee gravity: although imposing a vanishing potential fixes the functional form of the bumblebee vector $B_{\mu}$ and reduces the resulting gravitational field equations to second order, the theory remains intrinsically a higher-derivative gravity. In the limit $\ell\to 0$, the solution reduces to the standard dyonic RN black hole with a generic topological horizon. Alternatively, in the absence of both electric and magnetic charges with $k=1$, the solution reduces to a Schwarzschild-like black hole, which has been extensively discussed in Ref. [16]. The Kretschmann scalar, i.e., the square of the Riemann tensor, can be easily calculated as
@@
which exhibits the similar singularity structure as the dyonic RN case, but clearly deviates from the corresponding Kretschmann invariant when $\ell\neq 0$. In particular, several terms contain $\ell$ as an overall prefactor in the numerator, with no counterparts in GR. These contributions vanish when $\ell=0$, confirming that they arise solely due to Lorentz-symmetry breaking.
@@ III.2 The difference from [18]
$h$ $=$ $k-\frac{m}{r}+\frac{(1+\ell)q^{2}}{2\left(2+\ell\right)r^{2}}\,,\quad f=\frac{h}{1+\ell}\,$,
$\phi$ $=$ $\frac{\sqrt{\ell+1}q}{r}\,,\quad b_{r}=\frac{b}{\sqrt{f}}\,.$ (31)
@@
$h$ $=$ $k-\frac{m}{r}+\frac{(1+\ell)q^{2}}{2\left(2+\ell\right)r^{2}}+\frac{(\ell+1)(3\ell+2)p^{2}}{2(\ell+2)^{2}r^{2}}\,$,
$\gamma_{2}$ $=$ $-\frac{2\gamma(\ell+1)}{(\ell+2)^{2}}\frac{p^{2}}{q^{2}}\,.$ (33)
@@
$A_{{{\scriptscriptstyle(1)}}}$ $=$ $\phi(r)dt+pqud\varphi\,$,
$h$ $=$ $k-\frac{m}{r}+\frac{(1+\ell)q^{2}}{2\left(2+\ell\right)r^{2}}+\frac{(\ell+1)(3\ell+2)p^{2}q^{2}}{2(\ell+2)^{2}r^{2}}\,$, (35)
@@ IV Thermodynamics
In the limit $\ell\to 0$, the temperature reduces to the GR result. Next, in order to obtain the electric charge $Q_{e}$, it is convenient to introduce an antisymmetric tensor ${\cal F}^{\mu\nu}$
@@
It is evident that the integration constant $m$ is proportional to the Komar mass. One may further verify that the same integration constant is also proportional to the Arnowitt-Deser-Misner mass. The Wald entropy formula [94, 95] reads
@@ IV.1 Wald formalism
A direct consequence of the Wald formalism is the Wald entropy formula, which extends the Bekenstein-Hawking entropy and possesses remarkable universality. In higher-derivative gravity theories, black hole entropy is no longer proportional to the horizon area. The Wald entropy formula, however, correctly accounts for these corrections [106, 107, 108] and is thus widely used in many black hole–related effective theories [113, 104, 105, 109, 110, 111, 112]. Thus far, the Wald formalism and the Wald entropy formula have been applied to a broad range of extended gravity theories, including tensor-scalar theories [114, 115, 102, 103], higher-derivative graivty theories coupled maxwell field [119, 120], Maxwell theory and its extensions [116, 117, 118, 116, 106, 99, 100, 121], as well as the neutral bumblebee theory [49, 19].
In fact, the Wald entropy formula is simply a corollary of the Wald formalism and can be obtained from it only under specific conditions. Indeed, previous studies of Horndeski theory and bumblebee theory have demonstrated that the Wald entropy formula fails to provide a black hole entropy consistent with the first law of thermodynamics [49, 19, 102, 103]. More precisely, the dependence of the Noether charge on the Killing vector involves only $\xi^{\mu}$ and $\nabla^{\mu}\xi^{\nu}$. On the black hole horizon, if all fields are smooth, one has $\xi\to 0$ and $\nabla^{\mu}\xi^{\nu}\to K\epsilon^{\mu\nu}$, from which the Wald entropy formula follows. However, the $B_{\mu}$ field (16, 17) diverges on the black hole horizon, rendering the Wald entropy formula inapplicable. Although this divergence—like that of $g_{rr}$—is purely a coordinate artifact, the complete thermodynamic quantities must nevertheless be derived directly from the original Wald formalism.
@@
In the limit $\ell\to 0$, the Noether mass (86) coincides with the Komar mass (47), reproducing the standard GR result as expected. At horizon, we have
@@
In the limit $\ell\to 0$, the entropy (88) reduce to the area entropy in GR. Thus the first law of black hole thermodynamics reads
@@
It is evident that, although the Lorentz-violating parameter $\ell$ affects nearly all thermodynamic quantities of the dyonic RN-like black hole—except for the electric potential and magnetic charge—all forms of the first law of black hole thermodynamics, Eqs. (89)–(91), remain consistent with their GR counterparts. Finally, since all relevant physical quantities have now been determined, it is convenient to rewrite the black-hole solution directly in terms of these physical quantities rather than in terms of the integration constants. The metric, electricmagnetic field, and bumblebee field can be expressed as
$ds^{2}$ $=$ $\left[k-\frac{\kappa M}{rw_{2,k}\sqrt{1+\ell}}+\frac{\kappa^{2}Q_{e}^{2}(2+\ell)}{2w_{2,k}^{2}r^{2}}+\frac{2\kappa^{2}\left(1+\ell\right)Q_{m}^{2}}{w_{2,k}^{2}\left(2+3\ell\right)r^{2}}\right]dt^{2}$
$+\frac{1+\ell}{k-\frac{\kappa M}{rw_{2,k}\sqrt{1+\ell}}+\frac{\kappa^{2}Q_{e}^{2}(2+\ell)}{2w_{2,k}^{2}r^{2}}+\frac{2\kappa^{2}\left(1+\ell\right)Q_{m}^{2}}{w_{2,k}^{2}\left(2+3\ell\right)r^{2}}}dr^{2}+r^{2}\left(\frac{du^{2}}{1-ku^{2}}+(1-ku^{2})d\varphi^{2}\right)\,$,
$A$ $=$ $\frac{\kappa(2+\ell)Q_{e}}{rw_{2,k}\sqrt{1+\ell}}dt+\frac{2\kappa Q_{m}u}{w_{2,k}}d\varphi\,,\quad B=\frac{b}{\sqrt{k-\frac{\kappa M}{rw_{2,k}\sqrt{1+\ell}}+\frac{\kappa^{2}Q_{e}^{2}(2+\ell)}{2w_{2,k}^{2}r^{2}}+\frac{2\kappa^{2}\left(1+\ell\right)Q_{m}^{2}}{w_{2,k}^{2}\left(2+3\ell\right)r^{2}}}}\,$,
$\gamma_{1}$ $=$ $\frac{\gamma}{4(2+3\ell)}\,,\quad\gamma_{2}=-\frac{2\gamma(1+\ell)}{(2+\ell)(2+3\ell)}\,,\quad\ell=b^{2}\gamma\,.$ (92)
For the spherical topological case, i.e. $k=1$, and $w_{2,k}=4\pi$, the dyonic RN-like black hole possesses two horizons,
$r_{\pm}=\frac{M\pm\sqrt{M^{2}-2(\ell+1)\left((\ell+2)Q_{e}^{2}+\frac{4(\ell+1)Q_{m}^{2}}{3\ell+2}\right)}}{\sqrt{\ell+1}}\,$, (93)
where the mass $M$, electric charge $Q_{e}$, and magnetic charge $Q_{m}$ must satisfy the following condition:
$M^{2}\geq 2(\ell+1)\left((\ell+2)Q_{e}^{2}+\frac{4(\ell+1)Q_{m}^{2}}{3\ell+2}\right)\,.$ (94)
In the limit $Q_{e}\to 0$ or $Q_{m}\to 0$, the solution reduces to purely electric or purely magnetic black hole, respectively. In the limit $\ell\to 0$, the above solution smoothly reduces to the standard dyonic RN black hole in GR, as expected.
@@ V.1 Exact solution
$ds^{2}$ $=$ $-h(dt+2Nud\varphi)^{2}+\frac{dr^{2}}{f}+(r^{2}+N^{2})d\Omega_{2,k}^{2}\,$,
$A_{{{\scriptscriptstyle(1)}}}$ $=$ $\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Big[\phi(dt+2Nud\varphi)+pud\varphi\Big]+c_{e}dt\,$,
$B_{{{\scriptscriptstyle(1)}}}$ $=$ $b_{r}dr\,$, (95)
where $d\Omega_{2,k}$ is given by (15), $N$ is called NUT parameter, and
$h(r)$ $=$ $\frac{(r^{2}-N^{2})k}{r^{2}+N^{2}}-\frac{mr}{r^{2}+N^{2}}+\frac{p^{2}+q^{2}}{2(\ell+2)(r^{2}+N^{2})}\,$,
$\phi(r)$ $=$ $\frac{1}{2N}\left(q\sin X(r)+p\cos X(r)-p\right)\,$,
$b_{r}(r)$ $=$ $\frac{b}{\sqrt{f}}\,,\qquad f(r)=\frac{h(r)}{1+\ell}\,,\quad\ell=b^{2}\gamma\,$,
$X(r)$ $=$ $2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\,.$ (96)
For $p=q=0$, and $k=1$, the solution reduces to that given in Ref. [19]. For $N=0$, the solution reduces to that presented in Section III. For convenience, we redefine the magnetic charge parameter $p$, which differs from the one in (14) by a constant factor $\sqrt{(3\ell+2)/((\ell+1)(\ell+2))}$. The parameter $c_{e}$ is a gauge constant, which does not affect the solution. In previous RN-like solution, it was set to zero.
@@ V.2 Normal thermodynamic quantities
$T=\frac{1}{4\pi r_{h}\sqrt{\ell+1}}\left[1-\frac{p^{2}+q^{2}}{2(\ell+2)(r_{h}^{2}+N^{2})}\right]\,.$ (97)
Using the expression for the Maxwell field $A_{{{\scriptscriptstyle(1)}}}$ and its EOM, one can determine the electric potential $\Phi_{e}$ and the electric charge $Q_{e}$. Likewise, the magnetic charge $Q_{m}$ can be obtained via the Bianchi identity $dF_{{{\scriptscriptstyle(2)}}}=0$, and the magnetic potential $\Phi_{m}$ can be calculated by introducing the dual 1-form field $\mathcal{A}_{{{\scriptscriptstyle(1)}}}$ satisfying $\mathcal{F}_{{{\scriptscriptstyle(2)}}}=d\mathcal{A}_{{{\scriptscriptstyle(1)}}}$, which is given by
@@
with
$\mathcal{A}_{t}$ $=$ $-\frac{\sqrt{\ell+1}}{N(\ell+2)}(q\cos X-p\sin X)\,$, (99)
$\mathcal{A}_{\varphi}$ $=$ $-\frac{2u\sqrt{\ell+1}}{\ell+2}(q\cos X-p\sin X)\,.$ (100)
In the limit $N\to 0$, Eq. (98) reduces to the dyonic RN-like case (45), with the gauge parameter $c_{m}=0$ and $p\to p\sqrt{((\ell+1)(\ell+2))/(3\ell+2)}$. The electric and magnetic potentials $\Phi_{e}$ and $\Phi_{m}$ are given by
$\Phi_{e}$ $=$ $\frac{1}{2N}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Big[q\sin X(r_{h})+p\cos X(r_{h})-p\Big]\,$, (101)
$\Phi_{m}$ $=$ $\frac{\sqrt{\ell+1}}{N(\ell+2)}\Big[q-q\cos X(r_{h})+p\sin X(r_{h})\Big]\,.$ (102)
Similarly, in the $N\to 0$ limit, Eqs. (101)–(102) reduce to Eqs. (43) and (46).
Because of the Misner string, computing the electric charge requires careful treatment of the integration. We compute the electric charge $Q_{e}$ following the same procedure as in (42). Unlike in the static black hole case ($N=0$), the presence of the NUT parameter introduces a $dr\wedge d\varphi$ component in the expression for $\mathcal{F}_{{{\scriptscriptstyle(2)}}}$
$*\mathcal{F}_{{{\scriptscriptstyle(2)}}}$ $=$ $\mathcal{Q}_{tr}dt\wedge dr+\mathcal{Q}_{r\varphi}dr\wedge d\varphi+\mathcal{Q}_{u\varphi}du\wedge d\varphi\,$,
$\mathcal{Q}_{r\varphi}$ $=$ $-2Nu\mathcal{Q}_{tr}\,$,
$\mathcal{Q}_{tr}$ $=$ $\frac{2(\ell+1)(p+2N\phi)}{\sqrt{(\ell+2)(3\ell+2)}(r^{2}+N^{2})}\,,\qquad\mathcal{Q}_{u\varphi}=\frac{2\sqrt{3\ell+2}(r^{2}+N^{2})\phi^{\prime}}{(\ell+2)^{3/2}}\,.$ (103)
@@
$Q_{e}$ $=$ $-\frac{1}{2\kappa}\int d\varphi\Bigg[\int_{-1}^{1}\mathcal{Q}_{u\varphi}(r\to\infty,u)du-\int_{r_{h}}^{\infty}\mathcal{Q}_{r\varphi}(r,u)\Big|_{u=-1}^{u=1}dr\Bigg]$ (104)
$=$ $-\frac{w_{2}}{4\kappa}\mathcal{A}_{\varphi}|_{u=-1}^{u=1}=\frac{w_{2}\sqrt{\ell+1}}{\kappa(\ell+2)}\big[q\cos X(r_{h})-p\sin X(r_{h})\big]\,$,
can be determined. Here, the charge integral is evaluated over $S^{2}$ ($w_{2}=4\pi$); therefore, the $dt\wedge dr$ component does not contribute to it. In the next subsection, it will be shown that this component corresponds to the charge induced by the NUT parameter.
As in the case of the electric charge, we compute the magnetic charge $Q_{m}$ using the procedure described in (44). Here
$F_{{{\scriptscriptstyle(2)}}}$ $=$ $\mathcal{P}_{tr}dt\wedge dr+\mathcal{P}_{r\varphi}dr\wedge d\varphi+\mathcal{P}_{u\varphi}du\wedge d\varphi\,$,
$\mathcal{P}_{r\varphi}$ $=$ $-2Nu\mathcal{P}_{tr}\,$,
$\mathcal{P}_{tr}$ $=$ $-\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\phi^{\prime}\,,\qquad\mathcal{P}_{u\varphi}=\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\big(p+2N\phi\big)\,.$ (105)
We continue to employ the same integration method as for the electric charge, allowing the magnetic charge to be computed straightforwardly
$Q_{m}$ $=$ $\frac{1}{2\kappa}\int d\varphi\Bigg[\int_{-1}^{1}\mathcal{P}_{u\varphi}(r\to\infty,u)du-\int_{r_{h}}^{\infty}\mathcal{P}_{r\varphi}(r,u)\Big|_{u=-1}^{u=1}dr\Bigg]$ (106)
$=$ $\frac{w_{2}}{4\kappa}A_{\varphi}|_{u=-1}^{u=1}=\frac{w_{2}}{2\kappa}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Big[p\cos X(r_{h})+q\sin X(r_{h})\Big]\,.$